OpenStudy (anonymous):

Use the rule of logs to solve for x. Lnx^2=2ln4-4ln2. I need help please.

jhonyy9 (jhonyy9):

do you know the rules of logarithms ? log a + log b = ? log a - log b = ? log a^x = ?

OpenStudy (anonymous):

A×b , a/b , x

jhonyy9 (jhonyy9):

the 3rd how you have got it ?

jhonyy9 (jhonyy9):

so little help than you know that log a^x = xlog a can you us it in case of your exercise ?

jhonyy9 (jhonyy9):

than you know that 4 = 2^2 how can you rewriting the first term from the right side ?

jhonyy9 (jhonyy9):

so than you dont like cooperate nothing bye